# Boosting Graph Pooling with Persistent Homology

### v2

@@ 2 Related work
Persistent homology in GNNs. PH is a technique to calculate topological features of structured data, and many approaches have been proposed to use PH in graph machine learning due to the high expressiveness of topological features on graphs [17]. Since non-isomorphic graphs may exhibit different topological features, the combination of PH and the Weisfeiler-Lehman (WL) algorithm leads to stronger expressive power [38]. This encourages further exploration on equipping GNNs with topological features. [49] propose that message passing in GNNs can be effectively reweighted using topological features. [16] and [18] provide theoretical and practical insights that filtraions in PH can be purely learnable, enabling flexible usage of topological features in GNNs. However, existing methods tend to view PH merely as a tool for providing supplementary information to GNNs, resulting in unsatisfactory improvements and limited interpretability.
@@ 3 Background
$\mathbf{X}^{(k)}=\mathrm{M}\left(\mathbf{A},\mathbf{X}^{(k-1)};\mathbf{\theta}^{(k)}\right)$ (1)
@@
$\mathbf{A}^{(l)}=\mathbf{S}^{(l)\top}\mathbf{A}^{(l-1)}\mathbf{S}^{(l)}$ (3)
@@
These approaches differ from each other in the way to produce $\mathbf{S}$, which is used to inject a bias in the formation of clusters. In our work, we select three GP methods, i.e., DiffPool [47], MinCutPool [2], and DMoNPool [32], to cope with. Details of these methods are in Appendix A.
@@
Simplicial complexes stand as the focal point within the realm of algebraic topology. An assembly of simplices of certain dimensions constitutes a simplicial complex denoted as $K$. A graph can be seen as a low-dimensional simplicial complex that only contains 0-simplices (vertices) and 1-simplices (edges) [18]. The simplest kind of topological features describing graphs are Betti numbers, formally denoted as $\beta_{0}$ for the number of connected components and $\beta_{1}$ for the number of cycles.
Despite the limited expressive power of these two numbers, it can be improved by evaluating them alongside a filtration. Filtrations are scalar-valued functions of the form $f:V\cup E\rightarrow\mathbb{R}$ , which assigns each vertex and edge a value. Changes in the Betti numbers, named as persistent Betti numbers, can subsequently be monitored throughout the progress of the filtration: by considering a threshold $(a\in\mathbb{R})$, we can analyze the subgraph originating from the pre-image of $((-\infty,a])$ of $f$, denoted as $(f^{-1}((-\infty,a]))$. The image of $f$ leads to a finite set of values $a_{1}<\cdots<a_{n}$ and generates a sequence of nested subgraphs of the form $\emptyset\subseteq\mathcal{G}_{0}\subseteq\ldots\mathcal{G}_{k}\ldots\subseteq\mathcal{G}_{n}=\mathcal{G}$, where $\mathcal{G}_{k}=(V_{k},E_{k})$ is a subgraph of $\mathcal{G}$ with $V_{k}:=\left\{v\in V\mid f\left(\mathbf{x}_{v}\right)\leq a_{k}\right\}$ and $E_{k}:=\left\{(v,w)\in E\mid\max\left\{f\left(x_{v}\right),f\left(x_{w}\right)\right\}\leq a_{k}\right\}$. This process is also known as persistent homology (denoted as $ph(\cdot)$) on graphs. Typically, persistent Betti numbers are summarized in a persistence diagram (PD) as $\mathrm{ph}(\mathcal{G},f)=\{\mathcal{D}_{0},\mathcal{D}_{1},...\}$, which is made up of tuples $(a_{i},a_{j})\in\mathbb{R}^{2}$, with $a_{i}$ and $a_{j}$ representing the creation and destruction of a topological feature respectively (see Fig. 1(c)). The absolute difference in function values $|a_{j}-a_{i}|$ is called the persistence of a topological feature, where high persistence corresponds to features of the function, while low persistence is typically considered as noise [18, 37].
@@ 4.2 Topology-Invariant Pooling
The core of PH is the notion of filtration, which presents a challenging task to choose the right filtration. As the coarsened graph evolves in each training step, integrating PH into GP demands multiple computations of filtrations. To address this, we place recently proposed learnable filtration (LF) functions [16] to incorporate PH information for flexibility and efficiency. LF relies on node features and graph topology, which are readily available in GP. Consequently, LF can be seamlessly integrated into GP with minimal computational overhead. Specifically, we employ an MLP network $\Phi(\cdot)$ as the filtration function together with $\mathrm{sigmoid}(\cdot)$ to map node features $\mathbf{X}\in\mathbb{R}^{n\times d}$ into $n$ scalar values. Recently, an increasing amount of attention has been devoted to cycles [4, 48, 20] due to their significant relevance to downstream tasks in various domains such as biology [25], chemistry [33], and social network analysis [22]. Recognizing that cycles offer an intuitive representation of graph structure, we focus on the one-dimensional PD associated with cycles. For those edges do not form cycles, their creation and destruction are the same, leading to zero persistence. Following the standard way in GP (Eq. 2 3 4), we additionally propose the subsequent modules to inject PH into GP at both feature and topology levels.
@@
One major limitation of utilizing LF proposed in [16] is that the computation process is unaware of edge weights, i.e. edges with non-zero weights will be treated equally, so PH cannot directly extract meaningful topology from $\mathbf{A}^{(l)}$. Besides, rethinking GP in Eq. 3, the coarsened adjacency matrix has limited expressive power for two reasons. First, although $\mathbf{S}^{(l)}$ is a soft assignment matrix obtained by $\mathrm{softmax}(\cdot)$, each element still has nonzero values, i.e. $\mathbf{A}^{(l)}$ is always densely connected. Second, the edge weights may span a wide range by multiplication (see Appendix D for empirical evidence). These drawbacks hinder the stability and generalization power of the subsequent message passing layers [13]. None of the existing GP methods can handle these problems properly.
@@
$\mathbf{A}^{\prime(l)}=\mathrm{resample}\left(\frac{\mathbf{A}^{(l)}-\min(\mathbf{A}^{(l)})}{\max(\mathbf{A}^{(l)})-\min(\mathbf{A}^{(l)})}\right)$ (5)
@@
$\mathcal{D}_{1}$ $=\mathrm{ph}(\mathbf{A}^{\prime(l)},\mathrm{sigmoid}(\Phi(\mathbf{X}^{(l)})))$ (6)
$\mathbf{A}^{(l)}$ $=\mathbf{A}^{\prime(l)}\odot\mathrm{to\_dense}(\mathcal{D}_{1}[1]-\mathcal{D}_{1}[0])$
where $\odot$ is the Hadamard product, $\mathrm{to\_dense()}$ means transforming sparse representations in terms of edges to dense matrix representations, $\mathcal{D}_{1}[i]$ is the $i$-th value in each tuple of $\mathcal{D}_{1}$, and we denote the updated adjacency matrix after persistence injection still as $\mathbf{A}^{(l)}$ for notation consistency. Persistence injection can actually be regarded as a reweighting process. Since the filtration values are within $[0,1]$, $\mathbf{A}^{(l)}$ after persistence injection is guaranteed to have edge weights in the range of $[0,1]$ and is passed to the next pooling layer.
@@
The aforementioned mechanism can explicitly inject topological information into graphs, but it relies on the condition that the coarsened graph retains certain essential sub-topology. To this end, we propose an additional loss function to guide the GP process. Intuitively, the coarsened graph should exhibit similarity to the original graph in terms of topology. Since the computation of PH is differentiable, one possible approach is to directly minimize the differences between the PDs of the original graph and the coarsened graph. However, this implementation would require computing the Wasserstein distance between two PDs through optimal transport [45], which is intractable in training due to its complexity. Considering that our objective is to estimate the difference, we instead propose vectorizing the PDs and minimizing their high-order statistical features [34]. Specifically, we use several transformations (denoted as $\mathrm{transform(\cdot)}$) and concatenate the output, including triangle point transformation, Gaussian point transformation and line point transformation introduced in [6] to convert the tuples in PD into vector $\mathbf{h}_{t}$ $(t\in[1,m])$. We calculate the mean vector $\mathbf{\mu}$ as well as the second-order statistics as the standard deviation vector $\mathbf{\sigma}$ as:
$\mathbf{h}_{t}=\mathrm{transform(\mathcal{D}_{1})}$ (7)
@@
$\mathcal{L}_{topo}=\frac{1}{Ld}\sum_{l=1}^{L}\sum_{i=1}^{d}\left(\left(\mu_{i}^{(l)}\|\sigma_{i}^{(l)}\right)-\left(\mu_{i}^{(0)}\|\sigma_{i}^{(0)}\right)\right)^{2}$ (8)
@@ 4.3 Analysis
Theorem 4.1.
The 1-dimensional topological features computed by PH is sufficient enough to be at least as expressive as 1-WL in terms of distinguishing non-isomorphic graphs with self-loops, i.e. if the 1-WL label sequences for two graphs $\mathcal{G}$ and $\mathcal{G}^{\prime}$ diverge, there exists an injective filtration $f$ such that the corresponding 1-dimensional persistence diagrams $\mathcal{D}_{1}$ and $\mathcal{D}^{\prime}_{1}$ are not equal.
This result demonstrates that the 1-dimensional topological features contain sufficient information to potentially perform at least as well as 1-WL when it comes to distinguishing non-isomorphic graphs. We can then obtain the concluding remark that TIP is more expressive than other dense pooling methods by showing that there are pairs of graphs that cannot be distinguished by 1-WL but can be distinguished by TIP.
Proposition 1. TIP is invariant under isomorphism.
@@
Complexity. PH can be efficiently computed for dimensions 0 and 1, with a worst-case time complexity of $O(m\alpha(m))$, where $m$ represents the number of sorted edges in a graph. Here, $\alpha(\cdot)$ represents the inverse Ackermann function, which is extremely slow-growing and can essentially be considered as a constant for practical purposes. Therefore, the primary factor that affects the calculation of PH is the complexity of sorting all the edges, which is $O(m\log m)$. Our resampling and persistence encoding mechanism ensures that the coarsened graphs are sparse rather than dense, making our approach both efficient and scalable. We provide running time comparisons in Appendix E.2, which indicates that the inclusion of TIP does not impose a significant computational burden.
@@ 5 Experiments
Q1. Is PH capable of preserving topological information during pooling? Q2. How does PH affect graph pooling in preserving task-specific information?
@@ 5.1 Experimental Setup
To evaluate the capabilities of our model across diverse domains, we assess its performance on a variety of graph datasets commonly used in graph classification tasks. We select several benchmarks from TU datasets [30] and OGB datasets [19]. Specifically, we adopt molecular datasets NCI1, NCI109, and OGBG-MOLHIV, bioinformatics datasets ENZYMES, PROTEINS, and DD, as well as social network datasets IMDB-BINARY and IMDB-MULTI. Furthermore, to investigate the topology-preserving ability of our method, we conduct experiments on several highly structured datasets (ring, torus, grid2d) obtained from the PyGSP library. Appendix B.2 provides detailed statistics of the datasets.
@@ 5.2 Preserving Topological Structure
In this experiment, we study Q1 about the ability of PH to preserve topological structure during pooling. Specifically, we assess the topological similarity between the original and coarsened graphs $\mathcal{G}$ and $\mathcal{G^{\prime}}$, by comparing the Wasserstein distance associated with their respective PDs $\mathcal{D}_{1}$ and $\mathcal{D^{\prime}}_{1}$. This evaluation criterion is widely used to compare the topological similarity of graphs [45, 39]. We utilize Forman curvature on each edge of the graph as the filtration, which incorporates edge weights and graph clusters to better capture the topological features of the coarsened graphs [40, 42]. We consider the 1-Wasserstein distance $W\left(\mathcal{D}_{1},\mathcal{D}^{\prime}_{1}\right)=\inf_{\gamma\in\Pi\left(\mathcal{D}_{1},\mathcal{D}^{\prime}_{1}\right)}\mathbb{E}_{(x,y)\sim\gamma}[\|x-y\|]$ as the evaluation metric, where $\Pi(\cdot)$ is the set of joint distributions $\gamma(x,y)$ whose marginals are $\mathcal{D}_{1}$ and $\mathcal{D}^{\prime}_{1}$, respectively. Note that we are not learning a new filtration but keep a fixed one. Rather, we use learnable filtrations in training to enhance flexibility, and solely optimize $L_{topo}$ as the main objective.
Table 1: Results to show the topology-preserving ability. Wasserstein distance (lower is better) is used to assess the topological similarity. A bold value indicates the overall winner.
@@
Table 2: Test accuracy (higher is better) of graph classification on benchmark datasets. A bold value indicates the overall winner. Gray background indicates that TIP outperforms the base GP.
@@
DMoNPool-TIP 79.68 $\pm$ 1.38 78.46 $\pm$ 1.50 45.84 $\pm$ 5.32 79.73 $\pm$ 3.66 81.46 $\pm$ 2.96 74.25 $\pm$ 3.06 54.23 $\pm$ 2.64 76.70 $\pm$ 0.62
5.3 Graph Classification

### v3

@@ 2 Related work
Persistent homology in GNNs. PH is a technique to calculate topological features of structured data, and many approaches have been proposed to use PH in graph machine learning due to the high expressiveness of topological features on graphs [18]. Since non-isomorphic graphs may exhibit different topological features, the combination of PH and the Weisfeiler-Lehman (WL) algorithm leads to stronger expressive power [40]. This encourages further exploration on equipping GNNs with topological features. [52] propose that message passing in GNNs can be effectively reweighted using topological features. [16] and [19] provide theoretical and practical insights that filtrations in PH can be purely learnable, enabling flexible usage of topological features in GNNs. However, existing methods tend to view PH merely as a tool for providing supplementary information to GNNs, resulting in unsatisfactory improvements and limited interpretability.
@@ 3 Background
$\mathbf{X}^{(k)}=\mathrm{M}\left(\mathbf{A},\mathbf{X}^{(k-1)};\mathbf{\theta}^{(k)}\right),$ (1)
@@
$\mathbf{A}^{(l)}=\mathbf{S}^{(l)\top}\mathbf{A}^{(l-1)}\mathbf{S}^{(l)},$ (3)
@@
These approaches differ from each other in the way to produce $\mathbf{S}$, which is used to inject a bias in the formation of clusters. In our work, we select three GP methods, i.e., DiffPool [50], MinCutPool [2], and DMoNPool [34], to cope with. Details of the pooling layers in these methods are summarized in Appendix A.
@@
A simplicial complex $K$ consists of a set of simplices of certain dimensions. Each simplex $\gamma\in K$ has a set of faces, and each face $\tau\in\gamma$ has to satisfy $\tau\in K$. An element $\gamma\in K$ with $|\gamma|=k+1$ is called a $k$-simplex, which we denote by writing $\mathrm{dim}\ \gamma=k$. Furthermore, if $k$ is maximal among all simplices in $K$, then $K$ is referred to as a $k$-dimensional simplicial complex. A graph can be seen as a low-dimensional simplicial complex that only contains 0-simplices (vertices) and 1-simplices (edges) [19]. The simplest kind of topological features describing graphs are Betti numbers, formally denoted as $\beta_{0}$ for the number of connected components and $\beta_{1}$ for the number of cycles.
Despite the limited expressive power of these two numbers, it can be improved by evaluating them alongside a filtration. Filtrations are scalar-valued functions of the form $f:V\cup E\rightarrow\mathbb{R}$. Changes in the Betti numbers, named as persistent Betti numbers, can subsequently be monitored throughout the progress of the filtration: by considering a threshold $(a\in\mathbb{R})$, we can analyze the subgraph originating from the pre-image of $((-\infty,a])$ of $f$, denoted as $(f^{-1}((-\infty,a]))$. The image of $f$ leads to a finite set of values $a_{1}<\cdots<a_{n}$ and generates a sequence of nested subgraphs of the form $\emptyset\subseteq\mathcal{G}_{0}\subseteq\ldots\mathcal{G}_{k}\ldots\subseteq\mathcal{G}_{n}=\mathcal{G}$, where $\mathcal{G}_{k}=(V_{k},E_{k})$ is a subgraph of $\mathcal{G}$ with $V_{k}:=\left\{v\in V\mid f\left(\mathbf{x}_{v}\right)\leq a_{k}\right\}$ and $E_{k}:=\left\{(v,w)\in E\mid\max\left\{f\left(x_{v}\right),f\left(x_{w}\right)\right\}\leq a_{k}\right\}$. This process is also known as persistent homology (denoted as $ph(\cdot)$) on graphs. Typically, persistent Betti numbers are summarized in a persistence diagram (PD) as $\mathrm{ph}(\mathcal{G},f)[i]=\mathcal{D}_{i}$, where $i\in[0,1,...]$ is the dimension of topological features. PD is made up of tuples $(a_{i},a_{j})\in\mathbb{R}^{2}$, with $a_{i}$ and $a_{j}$ representing the creation and destruction of a topological feature respectively (see Fig. 1(c)). The absolute difference in function values $|a_{j}-a_{i}|$ is called the persistence of a topological feature, where high persistence corresponds to features of the function, while low persistence is typically considered as noise [19, 39].
@@ 4.2 Topology-Invariant Pooling
The core of PH is the notion of filtration, the selection of which presents a challenging task. As the coarsened graph evolves in each training step, integrating PH into GP demands multiple computations of filtrations. To address this, we place recently proposed learnable filtration (LF) functions [16] to incorporate PH information for flexibility and efficiency. LF relies on node features and graph topology, which are readily available in GP. Consequently, LF can be seamlessly integrated into GP with minimal computational overhead. Specifically, we employ an MLP network $\Phi(\cdot)$ as the filtration function together with $\mathrm{sigmoid}(\cdot)$ to map node features $\mathbf{X}\in\mathbb{R}^{n\times d}$ into $n$ scalar values. Recently, an increasing amount of attention has been devoted to cycles [4, 51, 21] due to their significant relevance to downstream tasks in various domains such as biology [26], chemistry [35], and social network analysis [23]. Recognizing that cycles offer an intuitive representation of graph structure [31, 17], and preliminary experiments, shown in Appendix E.5, indicate that the additional inclusion of zero-dimensional topological features merely increases runtime, thus we instead focus on the one-dimensional PDs associated with cycles. For those edges do not form cycles, their creation and destruction are the same, leading to zero persistence. Following the standard way in GP (Eq. 2 3 4), we additionally propose the subsequent modules to inject PH into GP at both feature and topology levels.
@@
One major limitation of utilizing LF proposed in [16] is that the computation process is unaware of edge weights, i.e. edges with non-zero weights will be treated equally, so PH cannot directly extract meaningful topology from $\mathbf{A}^{(l)}$. Besides, rethinking GP in Eq. 3, the coarsened adjacency matrix has limited expressive power for two reasons. First, although $\mathbf{S}^{(l)}$ is a soft assignment matrix obtained by $\mathrm{softmax}(\cdot)$, each element still has nonzero values, i.e. $\mathbf{A}^{(l)}$ is always densely connected. Second, the edge weights may span a wide range by multiplication (refer to Appendix D for empirical evidence). These drawbacks hinder the stability and generalization power of the subsequent message passing layers [13]. None of the existing GP methods can handle these problems properly.
@@
$\mathbf{A}^{\prime(l)}=\mathrm{resample}\left(\frac{\mathbf{A}^{(l)}-\min(\mathbf{A}^{(l)})}{\max(\mathbf{A}^{(l)})-\min(\mathbf{A}^{(l)})}\right),$ (5)
@@
$\mathcal{\tilde{D}}_{1}$ $=\mathrm{ph}(\mathbf{A}^{\prime(l)},\mathrm{sigmoid}(\Phi(\mathbf{X}^{(l)})))[1]$ (6)
$\mathbf{A}^{(l)}$ $=\mathbf{A}^{\prime(l)}\odot\mathrm{to\_dense}(\mathcal{\tilde{D}}_{1}[1]-\mathcal{\tilde{D}}_{1}[0]),$
where $\odot$ is the Hadamard product, $\mathrm{to\_dense()}$ means transforming sparse representations in terms of edges to dense matrix representations, $\mathcal{\tilde{D}}_{1}$ is the augmented 1-dimensional PDs by placing the tuples correspond to self-loop edges on the diagonal part of original PDs $\mathcal{D}_{1}$, $\mathcal{\tilde{D}}_{1}[i]$ is the $i$-th value in each tuple of $\mathcal{\tilde{D}}_{1}$, and we denote the updated adjacency matrix after persistence injection still as $\mathbf{A}^{(l)}$ for notation consistency. Persistence injection can actually be regarded as a reweighting process. Since the filtration values are within $[0,1]$, $\mathbf{A}^{(l)}$ after persistence injection is guaranteed to have edge weights in the range of $[0,1]$ and is passed to the next pooling layer.
@@
The aforementioned mechanism can explicitly inject topological information into graphs, but it relies on the condition that the coarsened graph retains certain essential sub-topology. To this end, we propose an additional loss function to guide the GP process.
Intuitively, the coarsened graph should exhibit similarity to the original graph in terms of topology. Since the computation of PH is differentiable, one possible approach is to directly minimize the differences between the PDs of the original graph and the coarsened graph. However, this implementation would require computing the Wasserstein distance between two PDs through optimal transport [48], which is intractable in training due to its complexity. Considering that our objective is to estimate the difference, we instead propose vectorizing the PDs and minimizing their high-order statistical features [36]. Specifically, we use several transformations (denoted as $\mathrm{transform(\cdot)}$) and concatenate the output, including triangle point transformation, Gaussian point transformation and line point transformation introduced in [6] to convert the tuples in PD into vector $\mathbf{h}_{t}$ $(t\in[1,m])$. We calculate the mean vector $\mathbf{\mu}$ as well as the second-order statistics as the standard deviation vector $\mathbf{\sigma}$ as:
$\mathbf{h}_{t}=\mathrm{transform(\mathcal{\tilde{D}}_{1})}$ (7)
@@
$\mathcal{L}_{topo}=\frac{1}{Ld}\sum_{l=1}^{L}\sum_{i=1}^{d}\left(\left(\mu_{i}^{(l)}\|\sigma_{i}^{(l)}\right)-\left(\mu_{i}^{(0)}\|\sigma_{i}^{(0)}\right)\right)^{2},$ (8)
@@ 4.3 Analysis
Theorem 1.
The self-loop augmented 1-dimensional topological features computed by PH is sufficient enough to be at least as expressive as 1-WL in terms of distinguishing non-isomorphic graphs with self-loops, i.e. if the 1-WL label sequences for two graphs $\mathcal{G}$ and $\mathcal{G}^{\prime}$ diverge, there exists an injective filtration $f$ such that the corresponding 1-dimensional persistence diagrams $\mathcal{\tilde{D}}_{1}$ and $\mathcal{\tilde{D}}^{\prime}_{1}$ are not equal.
Proof Sketch. We first assume the existence of a sequence of WL labels and show how to construct a filtration function $f$ from this. Consider nodes $u$ and $u^{\prime}$ are nodes with unique label count in $\mathcal{G}$ and $\mathcal{G^{\prime}}$, then our filtration is constructed such that their filtration values $f(u)$ and $f(u^{\prime})$ are unique and different. Consider all three cases: (1) $u$ and $u^{\prime}$ are both in cycles; (2) $u$ and $u^{\prime}$ are both not in cycles; (3) one of $u$ and $u^{\prime}$ is in cycles and the other is not. For all the cases, $f(u)$ and $f(u^{\prime})$ will be revealed in their respective persistence diagrams. Since $f(u)$ and $f(u^{\prime})$ are unique and different, we can use the augmented persistence diagrams to distinguish the two graphs.
This result demonstrates that the self-loop augmented 1-dimensional topological features contain sufficient information to potentially perform at least as well as 1-WL when it comes to distinguishing non-isomorphic graphs. We can then obtain the concluding remark that TIP is more expressive than other dense pooling methods by showing that there are pairs of graphs that cannot be distinguished by 1-WL but can be distinguished by TIP. Besides, our proposed simple yet effective self-loop augmentation eliminates the necessity of computing 0-dimensional topological features, thus reducing computational burdens.
Proposition 1.
TIP is invariant under isomorphism.
@@
Complexity. PH can be efficiently computed for dimensions 0 and 1, with a worst-case time complexity of $O(m\alpha(m))$, where $m$ represents the number of sorted edges in a graph. Here, $\alpha(\cdot)$ represents the inverse Ackermann function, which is extremely slow-growing and can essentially be considered as a constant for practical purposes. Therefore, the primary factor that affects the calculation of PH is the complexity of sorting all the edges, which is $O(m\log m)$. Our resampling and persistence injection mechanism ensures that the coarsened graphs are sparse rather than dense, making our approach both efficient and scalable. We provide running time comparisons in Appendix E.2, which indicates that the inclusion of TIP does not impose a significant computational burden.
@@ 5 Experiments
Q1. Is PH capable of preserving topological information during pooling?
Q2. How does PH affect graph pooling in preserving task-specific information?
@@ 5.1 Experimental Setup
To evaluate the capabilities of our model across diverse domains, we assess its performance on a variety of graph datasets commonly used in graph related tasks. We select several benchmarks from TU datasets [32], OGB datasets [20] and ZINC dataset [43]. Specifically, we adopt molecular datasets NCI1, NCI109, and OGBG-MOLHIV, bioinformatics datasets ENZYMES, PROTEINS, and DD, as well as social network datasets IMDB-BINARY and IMDB-MULTI. Furthermore, to investigate the topology-preserving ability of our method, we conduct experiments on several highly structured datasets (ring, torus, grid2d) obtained from the PyGSP library. Appendix B.2 provides detailed statistics of the datasets.
Evaluation.
@@ 5.2 Preserving Topological Structure
In this experiment, we study Q1 about the ability of PH to preserve topological structure during pooling. Specifically, we assess the topological similarity between the original and coarsened graphs $\mathcal{G}$ and $\mathcal{G^{\prime}}$, by comparing the Wasserstein distance associated with their respective PDs $\mathcal{\tilde{D}}_{1}$ and $\mathcal{\tilde{D}^{\prime}}_{1}$. This evaluation criterion is widely used to compare the topological similarity of graphs [48, 41]. We utilize Forman curvature on each edge of the graph as the filtration, which incorporates edge weights and graph clusters to better capture the topological features of the coarsened graphs [42, 45]. We consider the 1-Wasserstein distance $W\left(\mathcal{\tilde{D}}_{1},\mathcal{\tilde{D}}^{\prime}_{1}\right)=\inf_{\delta\in\Pi\left(\mathcal{\tilde{D}}_{1},\mathcal{\tilde{D}}^{\prime}_{1}\right)}\mathbb{E}_{(x,y)\sim\delta}[\|x-y\|]$ as the evaluation metric, where $\Pi(\cdot)$ is the set of joint distributions $\delta(x,y)$ whose marginals are $\mathcal{\tilde{D}}_{1}$ and $\mathcal{\tilde{D}}^{\prime}_{1}$, respectively. Note that we are not learning a new filtration but keep a fixed one. Rather, we use learnable filtrations in training to enhance flexibility, and solely optimize $L_{topo}$ as the main objective.
Table 1: Results to show the topology-preserving ability. Wasserstein distance ($\downarrow$) is used to assess the topological similarity. A bold value indicates the overall winner.
@@
Table 2: Test accuracy ($\uparrow$) of graph classification on benchmark datasets. A bold value indicates the overall winner. Gray background indicates that TIP outperforms the base GP.
@@
DMoNPool-TIP 79.68 $\pm$ 1.38 78.46 $\pm$ 1.50 45.84 $\pm$ 5.32 79.73 $\pm$ 3.66 81.46 $\pm$ 2.96 74.25 $\pm$ 2.93 54.23 $\pm$ 2.64 76.70 $\pm$ 0.62
5.3 Preserving Task-Specific Information